OpenStudy (anonymous):

A car completes a journey at an average speed of 40 m.p.h. At what speed must it travel on the return journey if the average speed for the complete journey (out and back) is 60 m.p.h?

OpenStudy (a_clan):

Average speed = Total Distance/ Total Time 60 = 2D / (t1 + t2) We know. t1= D/40 and t2 = D/S ,where S is the speed of the return journey Substitute t1 and t2 in first equation